A mixture of 40 liters of milk and water contains of 10% water. how much water must be added to make 20% of water in the new mixture?
Answer by jorel1380(3719)

40x.1=4 liters of pure water
4+w=.2(40+w)
4+w=8+.2w
.8w=4
w=4/.8=5
You need 5 liters of water to get 20% water in your mix
45x.2=9
4+5=9..